# Low-Complexity Monitoring and Compensation of Transceiver IQ Imbalance by Multi-dimensional Architecture for Dual-Polarization 16 Quadrature Amplitude Modulation

### v1

@@ 3 Experiment
We investigate the performance of the proposed scheme through experiments. The experimental setup is shown in Fig. 15. Two DFB lasers with a linewidth <100 kHz and a wavelength = 1551.12 nm are used as the signal carrier and local oscillator, respectively. A 92 GSa/s arbitrary waveform generator (AWG, Keysight M8196A) is used to generate the four tributaries of the 36 GBaud DP-Nyquist-16QAM signal with a ROF = 0.5, and the signal is resampled to 2.5 SPS. Then modulated onto the optical carrier using a dual-polarization optical IQ modulator(DPIQ, FTM7977) with a 3-dB bandwidth of 23 GHz. The optical signal is amplified by two EDFAs and transmitted over a 100-km standard single-mode fiber(SSMF). At the receiver side, the optical signal is demodulated by an integrated coherent receiver (ICR, CPRV1225A) with a 3-dB bandwidth of 25 GHZ. Finally, the signal was sampled by an oscilloscope (OSC, DSOZ594A) at a sample rate of 80 GHz. In TX DSP, the CD is compensated by a pre-chirp filter with a block size of 112, and the TX IQ imbalance is added digitally. In RX DSP, the signal is resampled to 2 SPS, and the RX IQ imbalance is added digitally first. The signal is input to the CRA after the Gram-Schmidt orthogonalization procedure (GSOP). Meanwhile, the signal is input to the RX skew monitor to calculate the RX skew. In Fig. 15, we set the receiver’s X polarization(RX X pol.) IQ skew and Y polarization (RX Y pol.) IQ skew at 6 ps respectively. To verify the performance of our proposed methods, the experiments were repeated several times.
@@
Fig. 16 displays the performance of our proposed methods with the receiver IQ imperfect. The Fig. 16(a) shows the accuracy of the receiver IQ skew estimation. In Fig. 16(a), the skew estimation is performed twenty times. According to Fig. 16(a), most skew calculation results are around 6 ps, with a minimum of 5.8 ps and a maximum of 6.2 ps. This means our proposed scheme is effective and the estimation error is within 0.2 ps. The Fig. 16(b) displays the relation between Q-factor and receiver IQ amp./pha. imbalance after receiver IQ skew compensation. According to Fig. 16(b), the Q-factor is unaffected by IQ amp./pha. imbalance due to the GSOP.
@@
Fig. 17 displays the performance of our proposed methods with the transmitter IQ imperfect after RX X/Y pol skew compensation. The Fig. 17(a) shows the scanning module scans the TX delay result. In Fig. 17(a), the TX delay scans vary from -2T/16 to 2T/16, and the MSEs are plotted as a function of TX delay. The TX Y pol skew is set to about 2.5 ps. The TX Y pol skew is set to about 3.5 ps. The scanning delay is completely performed twice, which is the first scanning and the second scanning. According to The Fig. 17(a), the MSE is lowest when the TX scanning delay is 1.74 ps, both at the first and second scanning. Thus, the rough TX skew estimation is 1.74 ps. In Fig. 17(b), the TX amp./pha. imbalance is studied. The TX X/Y pol amplitude changes from -0.1 to 0.1, and the phase imbalance is respectively set to -10, 0, and 10 degrees. According to Fig. 17(b), the Q-factor remains stable when the amplitude imbalance varies from -0.1 to 0.1 with the phase imbalance fixed. On the other hand, the highest mean Q-factor is 8.7 dB when the TX phase imbalance is 0 degrees, and the lowest mean Q-factor is 8.3 dB when the TX phase imbalance is -10 degrees. This means that the Q penalty is 0.4 dB when using our proposed methods in the phase imbalance range of -10 to 10 degrees.
@@
We also compare the performance of the 4×4 MIMO and C-R AEQ as shown in Fig. 18. Because the 4×4 MIMO only compensates for RX X/Y skew, for fair comparison, in Fig. 18, both TX and RX X/Y pol skew are compensated. In addition, the 4×4 MIMO or C-R AEQ is used for equalizing the same signal data at each comparison. Fig. 18(a) displays the relation between Q-factor and taps. We define the non-butterfly and butterfly filter tap as shown in Eq. 21 and Eq. 22.
Where is the number of taps per filter. In Fig. 18(a), the signal data is at the condition of TX X/Y pol amplitude imbalance -0.1 and TX X/Y pol phase imbalance -10 degrees with zero RX X/Y pol imbalance. The tap of 4×4 MIMO changes from 10 to 48, and C-R AEQ changes from 9 to 29 taps. The CV-MIMO tap is fixed at 6 in C-R AEQ. According to Fig. 18(a), the Q-factor penalty is about 1 dB when comparing the 4×4 MIMO 28 taps and C-R AEQ 17 taps, or comparing the 4×4 MIMO 45 taps and C-R AEQ 27 taps. The reason is that the C-R AEQ equalizes the signal and compensates for the TX amp./pha. imbalance at the same time. In Fig. 18(b), the signal data is at the condition that TX X/Y pol. amplitude imbalance of -0.1 to 0.1, with zero TX X/Y pol. phase imbalance and zero RX X/Y pol. amp./pha. imbalance. Fig. 18(b) shows the number of taps difference between C-R AEQ and 4×4 MIMO under the same Q-factor. On the left Y-axis, the number of taps is plotted as a function of TX amplitude imbalance, for 4×4 MIMO and C-R AEQ. On the right Y-axis, the Q-factors are plotted as a function of TX amplitude imbalance, for 4×4 MIMO and C-R AEQ. According to Fig. 18(b), the C-R AEQ maintains approximately 14 taps(6 taps CV-MIMO+15 taps RV-DD-LMS), with the Q-factor remaining at 8.2 dB. In contrast, even if the TX amp./pha. imbalance is 0, the 4×4 MIMO requires at least 28 taps to achieve the same Q-factor.
@@ 4 Complexity
In this part, the complexity of our proposed methods is studied. The complexity of our proposed methods is shown by the number of real multiplications via comparison with other techniques. The signal oversampling rate is 2×symbol rate.
@@ 4.2 IQ Skew Estimation Complexity
In RX IQ skew estimation complexity area, we select the technique that RX skew estimation based on Godard’s phase error detection(Godard’s PD)[8] to compare with our proposed RX skew estimation technique in complexity. The complexity calculation includes phase error detection, RX skew estimation, and RX skew compensation. According to refrence [8], the Godard’s PD requires real multiplications per symbol, where the is number of point FFT. In general, is set to 128. Thus, the complexity is 61 real multiplications per symbol. In contrast, the Gardner’s PD for skew compensation just requires 18 real multiplications per symbol, and the detail of complexity is that phase error detection, RX skew estimation, and RX skew compensation require 8,2 and 8 real multiplications per symbol, respectively. The complexity is reduced by more than 70. The reason is that the XI, XQ, YI, and YQ all need to perform FFTs in Godard’s PD, which significantly increases complexity. In rough TX IQ skew estimation complexity area, the TX skew can be adjusted by adjusting the transmitter output delay. Thus, the TX skew compensation has zero complexity.
@@ 4.3 AEQ Complexity
The AEQ complexity is calculated by comparing the C-R AEQ and 4×4 MIMO. According to Fig. 18(b), the C-R AEQ requires 6+15 taps with 4×4 MIMO requiring 28 taps when achieve the same Q-factor. According to reference[23], the butterfly filter total requires real multiplications per symbol, and the non-butterfly filter requires real multiplications per symbol, where is the number of taps per filter. Thus, the C-R AEQ requires 216 real multiplications. In contrast, the 4×4 MIMO requires 448 real multiplications. The complexity is reduced by more than 51.
@@ 5 Conclusion
We proposed a low-complexity IQ imbalance compensation architecture. First, the transceiver IQ skew is estimated and compensated for by Gardner’s PD and scanning delay. Then, the C-R AEQ consists of a 6-tap CV-MIMO and 15-tap RV-DD-LMS MIMO to equalize the signal and compensate for the TX phase/amplitude imbalance. The simulation and experimental results show that the RX skew estimation error is less than 0.2 ps, and the C-R AEQ can tolerate TX IQ phase/amplitude imbalance. In 100 km 36 GBaud DP-16QAM signal transmission, the complexity is reduced by about 51%.

### v2

@@ 3 Experiment
We investigate the performance of the proposed scheme through experiments. The experimental setup is shown in Fig. 16. Two DFB lasers with a linewidth <100 kHz and a wavelength = 1551.12 nm are used as the signal carrier and local oscillator, respectively. A 93.4 GSa/s arbitrary waveform generator (AWG, Keysight M8196A) is used to generate the four tributaries of the 36/50 GBaud DP-Nyquist-16QAM signal with a ROF = 0.5/0.2, and the signal is resampled to 2.5/1.85 SPS. Then the signal is modulated onto the optical carrier using a dual-polarization optical IQ modulator(DPIQ, FTM7977) with a 3-dB bandwidth of 23 GHz. The optical signal is amplified by two EDFAs and transmitted over a 100-km standard single-mode fiber(SSMF). At the receiver, the optical signal is demodulated by an integrated coherent receiver (ICR, CPRV1225A) with a 3-dB bandwidth of 29 GHz. Finally, the signal was sampled by an oscilloscope (OSC, DSOZ594A) at a sample rate of 80 GHz. In TX DSP, the CD is compensated by a pre-chirp filter with a block size of 112. The TX IQ skew/pha. imbalance is added digitally. The TX amp. imbalance is added by AWG. In RX DSP, the signal is resampled to 2 SPS, and the RX IQ imbalance is added digitally first. The signal is input to the CRA after the Gram-Schmidt orthogonalization procedure (GSOP). Meanwhile, the signal is input to the GSMA to calculate the RX skew. To verify the performance of our proposed methods, the experiments were repeated several times. The 36 GBaud signal is used to test the SDA and GSMA performance.
@@
Figure (17) characterizes the performance of GSMA with the receiver IQ imperfect. The RX skew is set to 6 ps. In Fig. (17)(a), the skew estimation is performed 20 times under each kind of IQ imbalance. As shown in Fig. (17)(a), most estimation results are around 6 ps, with a minimum of 5.8 ps and a maximum of 6.25 ps. This means the estimation error is less than 0.25 ps. Figure (17)(b) displays the relation between Q-factor and receiver IQ amp./pha. imbalance after receiver IQ skew compensation. The Q-factor is unaffected by IQ amp./pha. imbalance due to the GSOP.
@@
Figure (18) characterizes the performance of our proposed methods with the transmitter IQ imperfect after RX X/Y pol skew compensation. In Fig. (18)(a), the TX X/Y pol skew is set to 2.6/2.6 ps. The is plotted as a function of TX delay. According to Fig. (18)(a), the is minimized at TX scanning delays of 1.74 or 3.47 ps. We select 2 ps skew as the TX skew estimation result. In Fig. (18)(b), the TX amp./pha. imbalance is studied. The TX X/Y pol amplitude imbalance changes from -0.1 to 0.1, and the phase imbalance is set to -10, 0, and 10 deg respectively. Figure (18)(b) shows that the Q-factor penalty is 0.6 dB under a phase imbalance of 10 deg with the -0.1 amplitude imbalance.
@@
We compare the performance of the 4×4 MIMO and C-R AEQ as shown in Fig. (19). Because the 4×4 MIMO only compensates for RX X/Y IQ imbalance, for fair comparison, both TX and RX X/Y pol skew are compensated. Moreover, the same received signal is equalized using a 4×4 MIMO or C-R AEQ in each comparison. First, we define the non-butterfly and butterfly structure numbers of filter taps shown in Eq. 23 and Eq. 24.
Where is the number of taps per filter, is the number of taps in the non-butterfly filter, and is the number of taps in the butterfly filter. The CV-MIMO taps are fixed at 6 in C-R AEQ. In Fig.(19)(a) or (b), the TX pha. imbalance is -10 deg or 10 deg, respectively. According to Fig. 19(a) and (b), the Q-factor using C-R AEQ with 17 taps is higher than the Q-factor using 4×4 MIMO with 34 taps. In Fig.(19)(c), the C-R AEQ taps are reduced to 14, and the 4×4 MIMO taps are reduced to 28. It can be observed that the Q-factor penalty is less than 0.2 dB using C-R AEQ under zero TX amplitude imbalance.
We also evaluate the performance of C-R AEQ for a 50 GBaud 16QAM system, as shown in Fig.(20). In Fig.(20)(a) and (b), the C-R AEQ and the 4×4 MIMO number of taps are set to 23 and 46, respectively. In Fig.(20)(c), the C-R AEQ and the 4×4 MIMO number of taps are set to 20 and 40, respectively. As shown in Fig.(20)(a) and (b), the C-R AEQ can reduce the Q-factor penalty under TX IQ imbalance. The Fig.(20)(c) shows that, at zero IQ imbalance condition, the C-R AEQ needs 20 taps to achieve 8.1 dB Q-factor, and the 4×4 MIMO requires 40 taps to achieve the same Q-factor.
@@ 4 Complexity
In this part, the complexity of our proposed architecture is studied. The complexity is shown by the number of real multiplications via comparison with other techniques. The signal oversampling rate is twice the symbol rate.
@@ 4.2 IQ Skew Estimation Complexity
To assess the complexity of the GSMA, we compare GSMA with Godard’s PD-based RX skew estimation [10]. According to reference [10], Godard’s PD requires real multiplications per symbol, where is number of point FFT. Typically, is set to 128. Thus, the complexity is 61 real multiplications per symbol. In contrast, Gardner’s PD for skew compensation requires 34 real multiplications per symbol, and the detail of complexity is that phase error detection and RX skew compensation require 26 and 8 real multiplications per symbol, respectively. The complexity is reduced by more than 55. The reason is that the XI, XQ, YI, and YQ all need to perform FFTs in Godard’s PD, which significantly increases complexity.
Regarding the complexity of SDA, first, the scanning delay block can be implemented by adjusting the transmitter output delay. Second, even without SDA, updating the equalizer coefficients requires equalizer error. Therefore, the TX skew compensation needs approximately zero real multiplications.
@@ 4.3 AEQ Complexity
The AEQ complexity is calculated by comparing the C-R AEQ and 4×4 MIMO AEQ. According to Fig. 19(c), the C-R AEQ and 4×4 MIMO AEQ require 14 taps(6-tap CV-MIMO and 16-tap TMMA-RV-MIMO) and 28 taps, respectively, to achieve the same Q-factor. According to ref.[24], the butterfly filter requires real multiplications per symbol, and the non-butterfly filter requires real multiplications per symbol. Thus, the C-R AEQ requires 224 real multiplications . In contrast, the 4×4 MIMO requires 448 real multiplications. For the 50 Gbaud system, the C-R AEQ requires 320 real multiplications, and the 4×4 MIMO requires 640 real multiplications. In addition, the C-R AEQ can recover the carrier phase with 0 real multiplications, whereas the BPS/ML[25] needs 64 real multiplications. Therefore, the C-R AEQ complexity is reduced by 50. The numbers of real multiplications per symbol are compared in Table I.
@@ 5 Conclusion
We proposed a low-complexity multi-dimensional IQ imbalance compensation architecture. First, the transceiver IQ skew is estimated and compensated for by GSMA and SDA. Then, the C-R AEQ consists of a CV-MIMO and a TMMA-RV-MIMO equalizer to perform signal equalization, CPR and TX phase/amplitude imbalance compensation. For GSMA, the simulation and experimental results show that the RX skew estimation error is less than 0.25 ps. Compared to Godard’s PD, the complexity of GSMA is reduced by 55%. For SDA, the simulation and experimental results show that the TX skew estimation error is less than 0.9 ps. For C-R AEQ, it can tolerate 0.1 TX IQ amplitude imbalance and 5 degrees with 0.3 dB penalty in the 36 GBaud signal. In 100 km 36 GBaud DP-16QAM signal transmission, compared to 4×4 MIMO, the complexity of C-R AEQ is reduced by more than 50%. A key advantage of the proposed GSMA is its hardware reuse. The algorithm directly utilizes the existing clock recovery circuit (CRA), thereby eliminating the need for dedicated circuitry and avoiding additional chip area consumption. Both the SDA and C-R AEQ are low-complexity algorithms as well. Therefore, the architecture can be deployed in hardware-resource limit area.